### v3

# WritingBench: A Comprehensive Benchmark for Generative Writing

Abstract
Recent advancements in large language models (LLMs) have significantly enhanced text generation capabilities, yet evaluating their performance in generative writing remains a challenge. Existing benchmarks primarily focus on generic text generation or limited in writing tasks, failing to capture the diverse requirements of high-quality written contents across various domains. To bridge this gap, we present WritingBench, a comprehensive benchmark designed to evaluate LLMs across 6 core writing domains and 100 subdomains, encompassing creative, persuasive, informative, and technical writing. We further propose a query-dependent evaluation framework that empowers LLMs to dynamically generate instance-specific assessment criteria. This framework is complemented by a fine-tuned critic model for criteria-aware scoring, enabling evaluations in style, format and length. The framework’s validity is further demonstrated by its data curation capability, which enables 7B-parameter models to approach state-of-the-art (SOTA) performance. We open-source the benchmark, along with evaluation tools and modular framework components, to advance the development of LLMs in writing.
WritingBench: A Comprehensive Benchmark for Generative Writing
Yuning Wu1, Jiahao Mei1,3, Ming Yan1thanks: Corresponding authors, Chenliang Li1, Shaopeng Lai1, Yuran Ren2, Zijia Wang2, Ji Zhang1, Mengyue Wu3, Qin Jin2*, Fei Huang1 1Alibaba Group, 2Renmin University of China, 3Shanghai Jiao Tong University {yuningwu, ym119608}@alibaba-inc.com, qjin@ruc.edu.cn
1 Introduction
Figure 1: WritingBench query example with color-coded requirements. The three black-bordered categories highlight essential requirements analyzed in follow-up assessments. Red phrases correlate with gray-shaded writing support materials.
Figure 2: Construction pipeline of WritingBench queries. The refinement pool contains five writing requirements (three core competencies with black borders) and one expression type (purple). Checked strategies refine initial queries into multi-requirement prompts (color-coded text) with red phrases referencing gray materials. Implementation details in Section 3.1.
In recent years, LLMs DeepSeek-AI et al. (2025); Hurst et al. (2024) have revolutionized text generation, demonstrating impressive performance across diverse applications, from generating creative content Mirowski et al. (2023); Lei et al. (2024) and assisting in education Wang et al. (2024a); Li et al. (2024b) to enhancing professional workflows Shao et al. (2024); Li et al. (2024a). However, generative writing, which requires high levels of coherence, creativity, logical reasoning, and stylistic precision, poses a unique challenge that existing evaluation methods fail to address adequately.
Current evaluation benchmarks for generative writing suffer from two major limitations: 1) Limited scope and diversity in task formulation; and 2) Inadequate evaluation metrics for complex writing tasks. First, there is a significant lack of specialized benchmarks covering a broad range of writing tasks. Most existing writing-oriented benchmarks are restricted to single domains, such as fictions Karpinska et al. (2024); Gómez-Rodríguez and Williams (2023), and their task formulations tend to be simplistic—often relying on single-sentence queries Bai et al. (2024) or a small set of instruction templates Paech (2023); Que et al. (2024). Additionally, many benchmarks use homogeneous input materials Que et al. (2024); Karpinska et al. (2024), limiting their ability to accommodate the complex and customized requirements inherent in real-world writing scenarios. As a result, they fail to capture the diversity and intricacies of practical writing tasks (see Figure 1). Second, current automatic evaluation metrics lack the robustness needed for a comprehensive and nuanced assessment of writing quality. While LLM-based evaluation methods show promise in capturing semantic meanings Shao et al. (2024); Que et al. (2024); Bai et al. (2024), they typically rely on a narrow set of predefined criteria (e.g., fluency and coherence). As LLMs continue to evolve with increasingly sophisticated writing capabilities, these static evaluation criteria and frameworks are inadequate for assessing the complex, multi-dimensional nature of writing, including creativity, argumentation strength, and domain-specific adherence.
To address these challenges, we introduce WritingBench, a comprehensive benchmark and robust framework for evaluating general-purpose writing. Our approach begins with a carefully designed secondary domain categorization, grounded in real-world writing needs. We develop a four-stage query construction pipeline (illustrated in Figure 2), where LLMs first generate and diversify writing queries, followed by human-driven material collection and optimization. This process ensures a diverse set of writing tasks with broad domain coverage, varied requirements, and integration of heterogeneous source of materials. To enable a more nuanced evaluation, we propose a query-dependent evaluation framework that dynamically generates five instance-specific criteria using LLMs, which are then scored by a fine-tuned critic model. Finally, we integrate the framework to filter writing-specific data and train a small-scale model to verify its ability in identifying high-quality writing samples.
Our primary contributions are as follows:
$\bullet$ We present WritingBench, an open-source writing benchmark comprising 1,239 queries across 6 primary domains and 100 subdomains,featuring style, format and length requirements. WritingBench supports extended-context generation with input ranging from tens to thousands of words, addressing real-world diversity. It facilitates systematic evaluation to identify improvement areas and highlights the potential of chain-of-thought (CoT) processes in creative tasks.
$\bullet$ We propose a query-dependent evaluation framework that integrates instance-specific criteria generation with a criteria-aware scoring model. It achieves 83% human alignment, significantly surpassing static-criteria baselines (65%, 59%). The effectiveness is further evidenced by its data curation capability - models trained with framework-filtered data match SOTA performance.
$\bullet$ We publicly release WritingBench, including its evaluation protocols, criteria generation tools with an integrated critic model, and writing-enhanced models, to foster further research. Available at: https://github.com/X-PLUG/WritingBench.
2 Related Work
2.1 Writing Benchmarks
Existing evaluation benchmarks suffer from significant limitations in domain coverage and task granularity. For instance, EQ-Bench encompasses story-related quires using templated queries Paech (2023), while LongBench-Write incorporates length constraints in 120 queries Bai et al. (2024); however, they both lack hierarchical domain taxonomies and multi-dimensional requirement specifications (e.g., style and format). Furthermore, most current benchmarks rely on fixed instruction templates, short contexts or materials predominantly from a single source Que et al. (2024); Karpinska et al. (2024); Marco et al. (2024), rendering them insufficient for addressing the complexity of real-world data needs. In contrast, our proposed benchmark fills these gaps by introducing 1,239 free-form queries distributed across 6 primary domains and 100 subdomains, with potential controls over style, format, and length, paired with inputs ranging from tens to thousands of words.
Benchmark Num Domains Requirement Input Token Free Query-Form Diverse Material-Source
Primary Secondary Style Format Length Avg Max
EQ-Bench 241 1 / 130 213 /
LongBench-Write 120 7 / 87 684 /
HelloBench 647 5 38 1,210 7,766
WritingBench (Ours) 1,239 6 100 1,546 19,361
Table 1: Comparison of existing writing benchmarks.
2.2 Evaluation Methods
Using LLMs as evaluators has become a prevalent approach for evaluating the quality of generated responses. Typically, researchers pre-define a fixed set of evaluation dimensions applicable across all test instances Gómez-Rodríguez and Williams (2023); Marco et al. (2024). For example, SuperCLUE Xu et al. (2023) employs three dimensions, whereas LongBench-Write Bai et al. (2024) adopts six dimensions (e.g., relevance and accuracy). HelloBench Que et al. (2024) introduces task-specific dimensions, but the dimensions remain consistent across all queries of a given task. Although the LLM-as-a-judge approach enhances scalability, static evaluation dimensions often fail to accommodate the diversity of writing styles and specifications. To address this limitation, recent work Liang et al. (2024) trains a model to dynamically generate evaluation dimensions for individual queries. However, the dimensions remains confined to a small predefined set. In contrast, our query-dependent evaluation framework leverages LLMs to generate diverse and instance-specific criteria while fine-tuning a dedicated critic model to perform the evaluation.
2.3 Writing-Enhanced Models
Although existing LLMs demonstrate exceptional writing capabilities, researchers strive to further enhance their overall writing proficiency. Recent models, such as Weaver Wang et al. (2024b) and Suri Pham et al. (2024), have exhibited notable domain-specific strengths. For instance, Weaver benefits from over 200B parameter pretraining, supporting four distinct writing domains, while LongWriter Bai et al. (2024) specializes in length constraints. However, these models experience substantial performance degradation in cross-domain scenarios and multi-constraint tasks. In this work, leveraging the effectiveness of our evaluation framework, we introduce writing-enhanced models trained on high-quality data, achieving high performance across various tasks.
3 WritingBench
Figure 3: Domain categories in WritingBench. Inner represents the 6 primary domains and outer depicts 100 secondary subdomains (Sub = subdomains per category, Num = dataset entries).
Category Num
Avg
Token
Max
Token
Domain
Academic & Engineering 187 1,915 15,534
Finance & Business 238 1,762 19,361
Politics & Law 226 2,274 18,317
Literature & Arts 242 1,133 9,973
Education 151 1,173 10,737
Advertising & Marketing 195 886 6,504
Requirement
Style 395 1,404 18,197
Format 342 1,591 18,197
Length 214 1,226 14,097
Length
<1K 727 443 994
1K-3K 341 1,808 2,991
3K-5K 94 3,804 4,966
5K+ 77 8,042 19,361
Table 2: Data statistics for WritingBench categorized by domain, requirement, and length.
In this section, we will mainly introduce the construction process of our WritingBench, and the query-dependent evaluation framework with the accompanied critic model for criteria-aware evaluation. Additionally, we train a writing-enhanced model to demonstrate the effectiveness of our framework’s data curation capability.
3.1 Benchmark Construction
WritingBench is constructed following a delicate pipeline (see Figure 2) that combines model-generated query refinement and human annotation, ensuring diversity and real-world relevance. The two-phrase construction process is illustrated below.
3.1.1 Model-Augmented Query Generation
This phase leverages LLMs to generate an initial set of writing queries, which are enriched and diversified through systematic guidance, with material suggestions provided as needed.
Phase 1: Initial Query Generation
We begin by constructing a two-tiered domain pool grounded in real-world writing scenarios, consisting of 6 primary domains and 100 secondary subdomains (see Figure 3 and Appendix A for detailed descriptions). The selected domains are designed to capture both traditional and emerging user needs for AI-assisted writing, encompassing categories such as Academic & Engineering, Finance & Business, Politics & Law, Literature & Art, Education, Advertising & Marketing. Leveraging the domain and subdomain tags, we prompt two different LLMs (ChatGPT-4o-latest and Claude-3.5-Sonnet) to generate initial writing queries that simulate realistic user requests to maximize diversity (see Appendix B.1).
Phase 2: Query Diversification
To improve the diversity and practical applicability of queries, while taking into account real-world needs, we design a set of query diversification strategies (see Appendix B.2) inspired by Xu et al. (2024), include three core requirements and three auxiliary dimensions:
$\bullet$ Stylistic adjustments (e.g., “Use a friendly and simple tone that kids can easily understand.”)
$\bullet$ Format specifications (e.g., “Follow the IEEE conference template”)
$\bullet$ Length constraints (e.g., “Generate a 500-word executive summary”)
$\bullet$ Personalization (e.g., “From the viewpoint of an educator with two decades of experience.”)
$\bullet$ Content specificity (e.g., “Detail the 2023 Q3 financial metrics”)
$\bullet$ Expression (e.g., Modify the query expression to be shorter.)
The first three categories represent typical writing capabilities, which we will evaluate through specialized assessments. While refining the queries, the LLM simultaneously provides material requirements (e.g., requesting financial reports as input for market analysis queries) (see Appendix B.3). This approach results in enriched queries accompanied by corresponding recommended reference materials.
Figure 4: Example of dynamically generating criteria for a writing query in WritingBench. Different background colors represent various types of requirements.
3.1.2 Human-in-the-Loop Refinement
This phase incorporates human expertise to verify model-generated queries and supplement material requirements, thereby ensuring their alignment with real-world applications and avoiding harmful content.
Phase 1: Material Collection
At this stage, we engage 30 trained annotators, each compensated at \$18 per hour and possessing specialized expertise proven through domain-knowledge tests. Their role is to collect necessary open-source materials (e.g., public financial statements or legal templates), guided by material requirements generated by LLMs. To minimize errors arising from parsing documents in diverse formats, the annotators carefully extract and verify the most pertinent text segments.
Phase 2: Expert Screening & Optimization
Next, five experts with LLM experience or industry expertise perform data screening. The experts conduct a delicate two-stage filtering process: (1) query adaptation: ambiguous or unrealistic queries are revised to better align with the provided materials and practical scenarios (e.g., adjusting a legal opinion query to reference specific clauses from the supplied statutes), (2) material pruning: redundant or irrelevant content is eliminated from the collected materials, ensuring that the context provided for writing tasks remained focused and relevant while guaranteeing the information is free from harmful content.
Finally, we construct WritingBench, a benchmark comprising 1,239 queries categorized using a two-tiered taxonomy. In comparison to existing writing benchmarks summarized in Table 1, WritingBench exhibits notable advantages in terms of the number of instances, domain diversity, requirement coverage, and variability in input lengths. The detailed statistical distribution of WritingBench is shown in Table 2.
3.2 Query-Dependent Evaluation Framework
Traditional LLM-as-a-judge evaluations typically rely on fixed evaluation criteria derived from general writing assessment conventions Shao et al. (2024); Bai et al. (2024); Que et al. (2024). However, such static criteria exhibit three critical limitations: (1) Domain exhaustiveness: fixed criteria are inadequate in adapting to specialized domains, failing to address unique characteristics found in areas like technical documentation or creative writing; (2) Requirement specificity: such criteria lack the flexibility to encompass specific requirements related to style, format, or length control; and (3) Material dependency: they are insufficient to verify whether responses appropriately utilize the provided reference materials, such as incorporating data points or maintaining narrative continuity.
To address these challenges, we propose a query-dependent evaluation framework that enables dynamic adaptation to diverse writing scenarios. Our approach comprises two phases:
Phase 1: Dynamic Criteria Generation
As illustrated in Figure 4, given a query $q$ in the WritingBench, the LLM is prompted to automatically generate a set of five evaluation criteria, $C_{q}=\{c_{1},\ldots,c_{5}\}$ (opting for five criteria reflects a common approach observed in many contemporary evaluation contexts Shao et al. (2024); Ma et al. (2024); Marco et al. (2024)), using a carefully designed instruction to ensure structural guidance during criteria specification (see Appendix B.4). Each criterion comprises three components: a concise name summarizing the criterion, an extended description elaborating on the evaluation focus, and detailed scoring rubrics, which provide fine-grained quality levels for the respective evaluation dimensions.
Phase 2: Rubric-based Scoring
For each criterion $c_{i}\in C_{q}$, the evaluator independently assigns a score on a 10-point scale to a response $r$, providing both a score and a justification. The final overall score is computed by averaging scores across all criteria. Detailed prompt is provided in Appendix B.5.
Models Avg Languages Domains Requirements
ZH EN D1 D2 D3 D4 D5 D6 R1 C R2 C R3 C
Proprietary LLMs
ChatGPT-4o-latest 8.16 8.3 8.1 8.1 8.1 8.2 8.1 8.4 8.1 8.3 8.7 8.2 8.9 8.2 8.3
o1-Preview 8.15 8.1 8.2 8.0 8.1 8.2 8.2 8.4 8.1 8.2 8.6 8.2 8.8 8.2 8.2
Claude-3.5-Sonnet 7.71 7.7 7.7 7.6 7.5 7.6 7.7 7.9 8.0 7.9 8.5 7.7 8.5 7.9 8.0
Gemini-1.5-Pro 7.78 7.8 7.7 7.7 7.5 7.8 7.9 8.0 7.9 7.9 8.6 7.9 8.8 7.9 8.0
Qwen-Max 8.37 8.4 8.3 8.3 8.3 8.4 8.4 8.5 8.4 8.5 8.7 8.4 9.0 8.4 8.5
Open-source LLMs
Deepseek-R1 8.55 8.7 8.5 8.5 8.5 8.6 8.6 8.7 8.6 8.7 8.9 8.6 9.0 8.6 8.7
Deepseek-V3 7.95 8.0 7.9 7.9 7.8 8.0 7.8 8.2 8.0 8.1 8.6 8.0 8.9 8.0 8.2
Mistral-Large-Instruct 7.64 7.6 7.7 7.7 7.6 7.8 7.3 7.9 7.6 7.7 8.2 7.7 8.7 7.7 7.9
Qwen-2.5-72B-Instruct 7.90 8.0 7.9 8.0 7.8 8.1 7.7 8.2 7.8 8.0 8.3 8.0 8.8 7.9 8.0
Qwen-2.5-7B-Instruct 7.43 7.3 7.5 7.7 7.4 7.6 6.9 7.8 7.3 7.5 7.9 7.6 8.6 7.4 7.5
Llama-3.3-70B-Instruct 7.01 6.7 7.3 7.0 6.9 7.0 6.8 7.3 7.3 7.1 7.8 7.1 8.2 7.0 7.2
Llama-3.1-8B-Instruct 6.35 5.7 6.9 6.6 6.4 6.1 6.0 6.7 6.6 6.4 7.0 6.4 7.6 6.3 6.4
Capability-enhanced LLMs
Suri 4.97 4.4 5.5 5.6 5.3 5.0 4.1 5.0 5.1 4.8 5.2 5.0 5.4 4.5 4.0
LongWriter 7.91 7.9 7.9 8.0 8.1 8.1 7.7 8.1 7.6 7.9 8.2 8.1 8.8 7.7 7.7
Qwen-2.5-7B-filtered 8.49 8.6 8.4 8.4 8.4 8.6 8.4 8.6 8.5 8.6 8.8 8.5 9.0 8.5 8.6
Llama-3.1-8B-filtered 8.49 8.6 8.4 8.5 8.4 8.6 8.4 8.6 8.5 8.6 8.8 8.5 8.9 8.5 8.5
Table 3: WritingBench performance of different LLMs across 6 domains and 3 writing requirements evaluated with our critic model (scale: 1-10). The six domains include: (D1) Academic & Engineering, (D2) Finance & Business, (D3) Politics & Law, (D4) Literature & Art, (D5) Education, and (D6) Advertising & Marketing. The three writing requirements assessed are: (R1) Style, (R2) Format, and (R3) Length. Here, “C" indicates category-specific scores.
3.3 Critic Model
To alleviate the computational overhead with LLM-based evaluation, we develop a dedicated critic model, $\mathcal{M}$, designed to implement our rubric-based scoring framework. Specifically, this model performs the mapping $\mathcal{M}_{c}:(q,r,C_{i})\mapsto[1,10]\times\mathcal{J}$, where the output consists of a numerical score and corresponding justification text, $\mathcal{J}$, both in accordance with the predefined evaluation rubric.
We fine-tune the critic model on a dataset comprising 50K instances, which are collected using LLMs in our experiments. The dataset encompasses diverse queries, evaluation criteria, and model responses to enhance the robustness of evaluation. The experiments presented in Section 4.3 validate the consistency of the critic model.
3.4 Evaluation-Guided Data Curation for Writing Enhancement
The query-dependent evaluation framework enables systematic data curation across diverse writing tasks through two-phase filtering. First, applying instance-specific criteria generation to initial 24K SFT samples. Subsequently, the critic model filters 50% samples, yielding 12K framework-curated samples with higher quality. Fine-tuning experiments are conducted using the llama-3.1-8b-instruct and qwen-2.5-7b-instruct models. Evaluations are performed on two writing benchmarks to validate the framework’s effectiveness in identifying high-quality writing samples.
4 Experiment
Figure 5: Scores of different models across different subdomains in WritingBench. Red indicates higher score and blue refers to lower score. The figures are the average score of each subdomain for different models.
4.1 Experiment Settings
This section outlines the settings employed in our experiments using the WritingBench framework. The evaluation protocol, and the model training configurations are as follows:
Evaluation Protocol: Uniform generation settings are applied to model responses on WritingBench, allowing up to 16,000 tokens. We set the temperature at 0.7, with a top-k of 20 and top-p of 0.8, ensuring a balance of creativity. The critic model uses the same parameters, with input length capped at 2,048 tokens for scoring stability.
Model Training: The critic model is fine-tuned on the Qwen-2.5-7B-Instruct base model, using the AdamW optimizer with a 7e-6 learning rate. It is trained on 50K SFT data for 3 epochs across 8xA100 GPUs (batch size 64, 8-step accumulation). The writing models, trained on both Qwen-2.5-7B-Instruct and Llama-3.1-8B-Instruct for 5 epochs on 32xA100 GPUs, achieving a total batch size of 128 with 4-step gradient accumulation.
4.2 Comparison between LLMs
We evaluate 16 LLMs: ChatGPT-4o-latest Hurst et al. (2024), o1-Preview Jaech et al. (2024), Claude-3.5-Sonnet Anthropic (2024), Gemini-1.5-Pro Reid et al. (2024), Qwen-Max Team (2025), Deepseek-R1 DeepSeek-AI et al. (2025), Deepseek-V3 DeepSeek-AI et al. (2024), Mistral-Large-Instruct MistralAI (2024), Qwen-2.5-72B-Instruct and Qwen-2.5-7B-Instruct Yang et al. (2024), Llama-3.3-70B-Instruct and Llama-3.1-8B-Instruct Dubey et al. (2024), Suri Pham et al. (2024), LongWriter Bai et al. (2024), and our writing model, Qwen-2.5-7B fine-tuned model and Llama-3.1-8B fine-tuned model, on WritingBench. Each query is assessed by 5 instance-specific criteria using a 10-point scale, with subdomain-specific subcategory heatmaps revealing detailed variations.
Key Insights from Domain Scores: From Table 3, we observe that Education (D5) always yields high performance among models, followed by Academic & Engineering (D1). Literature & Art (D5) is the lowest-scoring domain, showing notable variance. CoT-capable models like Deepseek-R1 and o1-preview outperform their non-CoT counterparts models, indicating CoT’s potential in handling narrative and creative content. Figure 5 identifies challenges in niche areas such as novel continuations, tender proposals, and white papers where models generally score lower. These tasks demand a higher level of knowledge, long-text generation capabilities, and adherence to contextual consistency, pinpointing areas for further enhancement.
Key Insights from Requirement Scores: Table 3 presents scores for three requirement subsets: R columns showing overall scores and C columns showing requirement-specific averages. For example, if a style query includes two relevant criteria (as confirmed by human annotation), the C column averages scores for those criteria only. Style (R1) and Format (R2) dimensions reveal model distinctions , with Deepseek-R1 and Qwen-Max often excels others. Nevertheless, length requirements (C of R3) remains challenging, especially in section-specific constraints (see Appendix 16 for examples) and extended text production (see Section 4.5).
The overall analysis of WritingBench experiment highlights: Deepseek-R1 consistently leads across both domain and requirement dimensions, showcasing its versatility and strong language capabilities. There is significant performance variance within creative content domains, with CoT models showing notable promise—a hypothesis further confirmed by our comparison (see Appendix C.1, where the model incorporating CoT mechanisms outperform the one without CoT on creative writing tasks). Additionally, we observe that advanced models often achieve higher specialized scores (C column) for the three common requirements than their overall scores (R column). The criteria outside these specialized sets tend to emphasize content aspects, such as integration with source materials and writing depth, highlighting the need to improve content quality. This analysis benchmarks existing model capabilities and underscores specific areas needing improvement for future development.
4.3 Human Consistency
To validate the alignment between automated evaluation and human judgment, we conducted human evaluation on 300 queries, covering all 100 subdomains. Five professionally trained annotators with linguistic backgrounds perform pairwise comparisons of model responses. For each query, two responses are randomly selected from different models. Annotators select the preferred response or declared equivalence based on the query’s requirements. We compare our dynamic query-dependent criteria against two baselines: static globally uniform criteria and static domain-specific customized criteria (static criteria are designed by domain experts) on two LLM judges: ChatGPT-4o-latest and Claude-3.5-Sonnet.
Evaluation Metric Judge Score
Static Global ChatGPT 69%
Static Domain-Specific ChatGPT 40%
Dynamic Query-Dependent ChatGPT 79%
Static Global Claude 65%
Static Domain-Specific Claude 59%
Dynamic Query-Dependent Claude 87%
Dynamic Query-Dependent Critic Model 83%
Table 4: Comparison of human consistency scores across different criteria generation methods. ChatGPT corresponds to ChatGPT-4o-latest, Claude corresponds to Claude-3.5-Sonnet.
As shown in Table 4, our dynamic query-dependent criteria achieve superior human alignment compared to static, both globally uniform or domain-specific customized criteria. Notably, domain-specific criteria underperform despite customization, since our queries involve highly diverse tasks and varied sources. These findings confirm that context-sensitive query-dependent evaluation better captures real-world writing complexity compared to conventional static approaches. Furthermore, the critic model attains 83% agreement, confirming its practical viability.
4.4 Ablation of Data Curation for Writing-Enhanced Models
Leveraging the query-dependent evaluation framework described in Section 3.2, we perform instance-specific criteria generation an initial 24K SFT dataset spanning various domains, including long-input queries enriched with supplemental materials. Our critic model then apply a query-dependent evaluation to filter this dataset down to a high-quality 12K subset. We conduct fine-tuning experiments using the Llama-3.1-8B-Instruct and Qwen-2.5-7B-Instruct models. As shown in 5, both models, trained on the filtered data, demonstrate significant performance improvements over their previous versions. Notably, they outperform models trained on the full dataset and even approached the capabilities of SOTA models. Further validation on an alternative benchmark, LongBench-Write (following the settings of quality evaluation in Bai et al. (2024)), confirm these findings, with our models surpassing baseline performance.
These results validate the robustness of our query-dependent evaluation strategy and the utility of our critic model for curating high-quality writing samples. This integrated approach enables smaller models to compete with, and even surpass, larger models across diverse writing tasks.
Models WritingBench Benchmark2
Deepseek-R1 8.55 4.79
Qwen-2.5-7B-Instruct 7.43 4.39
Llama-3.1-8B-Instruct 6.35 3.12
Qwen-2.5-7B-all 8.46 4.69
Qwen-2.5-7B-filtered 8.49 4.70
Llama-3.1-8B-all 8.45 4.65
Llama-3.1-8B-filtered 8.49 4.65
Table 5: Performance evaluation of our writing models on two benchmarks.’-filtered’ indicates models trained with filtered data, while ’-all’ refers to those trained with the full dataset.
4.5 Ablation of Length
We assess the performance of LLMs across varying input and output lengths, with statistical validity ensured by excluding intervals containing fewer than 5 samples.As demonstrated in Appendix 6, experiments reveal that most SOTA models generally maintain consistent performance regardless of input length variations, attributable to their advanced long-context comprehension capabilities.
However, analysis in Appendix 7 shows that most models exhibit inherent limitations in response generation length, typically producing outputs constrained to approximately 3,000 tokens. While output quality remains relatively stable within this range, smaller models suffer more performance degradation characterized by repetitive outputs. Notably, only two models, Qwen-Max and LongWriter, which incorporate architectural optimizations for long-form generation, effectively support extended response lengths. It underscores the importance of enhancing long-output capabilities and length optimization in writing.
5 Conclusion
In this paper, we introduce WritingBench, a comprehensive benchmark designed to evaluate LLMs in generative writing across diverse domains. It includes 1,239 queries spanning 6 primary domains and 100 subdomains, providing evaluation dimensions for style, format, and length requirements. Our query-dependent evaluation framework, supported by a critic model, achieves high human alignment. Evaluation efficiency is further demonstrated by compact models trained on curated data, achieving performance comparable to SOTA models. By making WritingBench and its resources publicly available, we aim to foster further research and advancements in LLM writing capabilities.
Limitations
This work faces three primary limitations that warrant consideration. First, both the writing model and critic model are primarily trained using conventional SFT approaches, omitting systematic exploration of enhanced optimization strategies. While we demonstrate the partial efficacy of the CoT mechanisms in creative domains, their potential remains unexplored compared to established successes in mathematical reasoning tasks.
Second, our evaluation framework exhibits diminished precision in handling complex multi-dimensional length requirements, including temporal sequencing constraints and section-specific word counts. This limitation underscores the necessity for enhanced scoring methodologies that integrate learned metrics with structured rule-based evaluations to better regulate output specifications.
Third, inherent challenges persist in obtaining reliable pairwise preference annotations for compositional tasks. Despite rigorous annotation protocols, human evaluators inevitably introduce subjective biases when assessing two fair-well responses, particularly regarding narrative preferences and contextual interpretations. While our consensus-building procedures mitigate some variability, absolute alignment with diverse user preferences remains theoretically unattainable.
Appendix A Benchmark Statistics
A.1 Overview of Six Primary Domains
  1. 1.
    Academic & Engineering: This domain encompasses the structured and formalized nature of academic writing workflows, focusing on clarity, precision, and adherence to rigorous standards. includes the creation of paper outlines, abstracts, literature reviews, experiment reports, and technical documents such as patents and test reports. The writing prioritizes logical argumentation, thorough analysis, and the integration of empirical evidence.
  2. 2.
    Finance & Business: Writing in this domain is analytical and strategic, aimed at informing decision-making and promoting corporate objectives. It includes a wide range of documentation such as contracts, market analyses, investment reports, strategic plans, and operational materials like product specifications and sales reports. The emphasis is on clarity and conciseness, with a focus on financial acumen and strategic insights.
  3. 3.
    Politics & Law: This domain demands an authoritative and formal tone, as it involves the composition of government documents, legal writings, and political communications. These materials require a careful balance between clarity and formality, often employing complex and structured language. The aim is to clearly convey policy positions, legal arguments, or political messages while strictly adhering to legal and procedural standards.
  4. 4.
    Literature & Art: This domain covers the creative and expressive realms of writing, including novels, poetry, scripts, artistic designs, and critiques of books and movies. Writers explore thematic and emotional depths, crafting works that connect with audiences on a human level. The language is rich and evocative, allowing for a personal exploration of ideas that engage and move the reader.
  5. 5.
    Education: This domain involves pedagogical materials and educational communication, including lesson plans, course designs, feedback, assignments, and institutional communications like admissions promotions and parent-teacher meeting scripts. The writing prioritizes clarity, accessibility, and instructional effectiveness, using an approachable tone to facilitate learning and engagement.
  6. 6.
    Advertising & Marketing: Writing in this domain is vibrant and persuasive, designed to captivate and influence target audiences across various digital platforms. It includes social media scripts, advertising copy, brand narratives, and multimedia campaign materials. The writing is dynamic and strategic, with a creative twist, necessitating a deep understanding of audience psychology and trend dynamics to ensure that the content is appealing and strategically effective.
A.2 Overview of 100 Secondary Subdomains
See Table 7, Table 8 and Table 9.
Appendix B Prompts
B.1 Initial Query Generation Prompt
Initial Query Generation Prompt Generate 10 different writing requests under {domain2} within the context of {domain1}. Ensure the requests are as detailed and specific as possible, and reflect realistic user tone and needs.
Please return in the following JSON format, and do not include anything outside of JSON:
   "Writing request 1",
   "Writing request 2",
Table 6: Initial query generation prompt introduced in Section 3.1.1.
Subdomain Description
Academic & Engineering
Paper Outline Hierarchical organization of research components and logical flow
Acknowledgments Formal recognition of institutional and individual support
Limitations Systematic identification of methodological constraints and scope boundaries
Defense Presentation Presentation supporting materials, such as slides
Research Proposal Investigation blueprint with validation road map
Technical Documentation Implementation specifications and system interface protocols
Experiments Parameterized validation framework with controlled variable analysis
Introduction Contextual foundation establishing research gaps and significance
Conclusion synthesize the main findings of the research or project
Test Report Evaluations of testing activities and performance
Contributions Novel aspects differentiating the work from prior research
Internship Report Chronological documentation of a practical work placement
Literature Review Critical gap analysis through scholarly works taxonomy
Defense Script Oral presentations and responses for research defense.
Abstract Summary of research objectives, methods, results, and significance
Engineering Report Technical analysis on tasks, methodologies, and outcomes
Patent Legal-technical specification of novel implementable claims
Finance & Business
Meeting Minutes Concise documentation of key discussion points, decisions, and action items
User Research Insight collection on user needs and behaviors to inform product or service design
Business Correspondence Formal communication with internal or external stakeholders for business purposes
Human Resource Management Strategies and processes for managing workforce effectively
Recruitment Strategies for attracting, selecting, and onboarding suitable candidates
Briefing Summarized information provided to stakeholders ahead of a task or meeting
Event Planning Coordinated organization of logistics and activities for event execution
Market Research Systematic collection and analysis about market and consumer
Market Analysis Evaluation of market trends, size, competitors, and dynamics
Risk Management Identification, assessment, and prioritization of risks with mitigation strategies
Sales Report Summary of sales activities, performance, and revenue figures over a given period
Pitch Deck Visual presentation designed to communicate business ideas or proposals to investors
Contract Legally binding agreement outlining the terms and conditions for business transactions
Tender Document Formal proposal request containing project specifications and bidding instructions
Investment Analysis Evaluation of financial investments to determine potential returns and risks
Product Proposal Detailed plan outlining the development, features, and potential of new products
Strategic Planning Business goal setting with actionable strategies for desired outcomes
Financial Reports Comprehensive statements reflecting the financial performance and status
Requirements Specification Documentation detailing functional and non-functional requirements for a project
Bid Proposal Formal offer to provide goods or services at a specified price, addressing client needs
Table 7: Subdomains in Academic & Engineering and Finance & Business.
Domain Description
Politics & Law
Legal Opinion Authoritative assessment and guidance on legal matters or questions
Government Speech Formal address delivered by government officials outlining policies or positions
Judgment Document Official written decision or order issued by a court
Legal Agreement Binding contract setting out terms and obligations between parties
Case Study In-depth analysis of a legal case for educational or professional purposes
Case Bulletin Summary and update on ongoing or concluded legal cases
Legal Consultation Professional advice provided on legal rights, responsibilities, or strategies
Regulatory Analysis Examination of rules and regulations affecting compliance and enforcement
Meeting Summary Brief overview of discussions, decisions, and outcomes from a meeting
Ideological Report Analysis or commentary on political or ideological trends and perspectives
Policy Interpretation Explanation or clarification for public or organizational guidance
Official Document Formal written record issued by government entities or officials
Legal Awareness Campaign Initiative to educate the public on legal rights and responsibilities
Defense Plea Formal written argument submitted by the defense in a legal proceeding
Party Membership Application Form and process for joining a political party
Policy Advocacy Efforts to influence or promote specific policy changes or implementations
Work Report Detailed account of activities, achievements, and challenges within a specific period
Deed Achievement Record highlighting significant accomplishments and contributions
Litigation Documents Legal filings and paperwork submitted in the course of a lawsuit
White Paper Authoritative report providing information or proposals on an issue
Literature & Art
Character Design Creation and development of detailed characters for stories or visual media
Greeting Message Friendly or formal introductory statement used for various occasions
Host Script Guided narration and dialogue for a presenter during an event or show
Novel Outline Structured plan for the plot, characters, and settings of a novel
Podcast Script Written content outlining the dialogue and segments for podcast episodes
Derivative Work Creative work based on or inspired by an existing piece
Reading Reflection Personal thoughts and analysis on a piece of literature
Video Script Script detailing dialogue and action for video content creation
Book Review Critical evaluation and summary of a book’s content and impact
Game Design Creation of game mechanics, stories, and interfaces for interactive entertainment
Lyric Writing Crafting of words for songs with rhyme and meter considerations
Brainstorm Rough ideas and notes generated during a creative thinking session
Plot Development Process of mapping out the storyline and narrative structure
Prose Written or spoken language in its ordinary form, without metrical structure
Screenplay Scripted blueprint for film or television with dialogue and directions
Novel Manuscript Complete text of a novel prepared for publication
Biography Detailed account of a person’s life experiences and achievements
Film/TV Review Analytical critique of a film or television show’s content and effectiveness
Poetry Artistic composition using rhythmic and metaphorical language
Fan Fiction Amateur stories written by enthusiasts featuring characters from existing media
Table 8: Subdomains in Politics & Law and Literature & Art.
Domain Description
Education
Training Reflection Personal assessment of training experiences and learned insights
Class Activity Planned exercises or tasks designed to engage students in learning
Parent-Teacher Meeting Formal discussion between educators and parents about student progress
Lesson Plan Structured outline of educational objectives and teaching methods for a class
Teaching Materials Resources used to aid in presenting information to students
Assignment Grading Evaluation and scoring of student work based on specific criteria
Curriculum Design Development of educational content, structure, and delivery methods
Educational Report Analysis or summary of educational outcomes and performance
Coursework Academic work assigned to students as part of a course
Evaluation Comments Feedback provided on student performance and areas of improvement
Educational Consulting Professional guidance on educational strategies and systems
Admissions Promotion Strategies and activities aimed at encouraging enrollment in educational institutions
Advertising & Marketing
Sales Letter Persuasive written communication intended to motivate potential buyers
Product Description Detailed overview of a product’s features, benefits, and uses
Social Media Content Engaging text, images, or videos crafted for online platforms
Multimedia Script Planned screenplay integrating various forms of media for marketing
Promotional Copy Compelling text written to boost interest and sales of products
Promotional Voiceover Recorded narration to accompany marketing visuals or ads
Travel Guide Informative content offering insights and tips for travelers
Brand Story Narrative that outlines the history, values, and mission of a brand
Personal Blog Individual commentary or stories shared in an informal online format
Marketing Commentary Analytical thoughts on marketing trends and strategies
Slogans Catchy and memorable phrases designed to convey brand identity
Table 9: Subdomains in Education and Advertising & Marketing.
B.2 Guidance Pool
Query Refinement Guidance Pool [
    "Add a requirement for generating specific lengths.",
    "Include format adherence requirements, such as writing according to a prescribed outline or outputting in a specific format.",
    "Add style requirements, like drafting a speech suitable for a particular occasion or adopting the style suitable for a specific audience or mimicking a particular tone.",
    "Incorporate user personalization needs, such as considering the user’s identity or integrating personal experiences.",
    "Include more specific content requirements, like details about a particular event or focusing on specific content.",
    "Express concisely in one sentence."
Table 10: Query refinement guidance pool introduced in Section B.5.
B.3 Query Refine Prompt
Query Refinement Prompt Please refine and enhance the original writing requirements in the context of generating content in {domain2} from {domain1} based on the provided guidance. Include as many details as possible and indicate whether additional writing materials are needed.
** Original Writing Requirements **
{query}
** Guidance for Modification **
{guidance}
** Output Requirements **
Return the result strictly in the following JSON format, with no additional content outside the JSON:
{     "query": "Modified writing requirements",
    "material": "Whether additional reference materials are needed to supplement the writing requirements. If needed, provide suggestions for the materials; if not needed, return"
Table 11: Query refinement prompt introduced in Section 3.1.1.
B.4 Criteria Generation Prompt
System Evaluation Prompt You are an expert evaluator with extensive experience in evaluating the response of a given query.
Table 12: Evaluation system prompt introduced in Section 3.2.
Criteria Generation Prompt Please generate five strict evaluation criteria for assessing the response given the following query. Each criterion should include the following fields: name, criteria_description, 1-2, 3-4, 5-6, 7-8, 9-10.
The criteria should be designed to emphasize detailed assessment and distinguish subtle differences in quality. Ensure that the criteria can discern issues such as relevance, coherence, depth, specificity, and adherence to the query context.
Do not include any additional text. Only output the criteria in the specified JSON format.
** Query **
{query}
** Output format **
   {"name": "first_criteria_name",
    "criteria_description": "Description for the first criteria, emphasizing detailed and critical assessment.",
    "1-2": "Low score description: Clearly deficient in this aspect, with significant issues.",
    "3-4": "Below average score description: Lacking in several important areas, with noticeable problems.",
    "5-6": "Average score description: Adequate but not exemplary, meets basic expectations with some minor issues.",
    "7-8": "Above average score description: Generally strong but with minor shortcomings.",
    "9-10": "High score description: Outstanding in this aspect, with no noticeable issues." },
Table 13: Criteria generation prompt introduced in Section 3.2.
B.5 Rubric-based Scoring Prompt
Scoring Prompt Evaluate the Response based on the Query and criteria provided.
** Criteria **
{criteria}
** Query **
{query}
** Response **
{response}
Provide your evaluation based on the criteria:
{criteria}
Provide reasons for each score, indicating where and why any strengths or deficiencies occur within the Response. Reference specific passages or elements from the text to support your justification.
Ensure that each reason is concrete, with explicit references to the text that align with the criteria requirements.
Scoring Range: Assign an integer score between 1 to 10
Output format:
Return the results in the following JSON format, Only output this JSON format and nothing else:
    "score": an integer score between 1 to 10,
    "reason": "Specific and detailed justification for the score using text elements."
Table 14: Scoring prompt introduced in Section 3.2.
Appendix C Experiment Results
C.1 Ablation of CoT in Creative content
To validate the impact of CoT reasoning on creative content generation, we conduct knowledge distillation experiments using DeepSeek-R1. Two variants of Qwen-2.5-32B-Instruct are developed: 1) Base model trained with CoT-formatted instructions, 2) ablated version (-w/o CoT) using direct response generation. These models are evaluated on the Literature & Art (D5) subset of WritingBench and EQBench, a specialized benchmark for fiction evaluation, to thoroughly examine their capabilities in generating creative content.
Models WritingBench-D4 EQBench
Deepseek-R1 8.55 84.99
Qwen-2.5-32B-Instruct 7.34 48.17
Qwen-2.5-32B-CoT 8.66 82.48
-w/o CoT 8.49 79.43
Table 15: Ablation of CoT in creative content on two benchmarks.
As shown in Table 15, the CoT-enhanced model achieves 8.66 (vs. 8.49) on WritingBench-D5 and 82.48 (vs. 79.43) on EQBench, showing CoT’s effectiveness in generating creative content.
C.2 Case Study
Query Create advertising slogans for a pet grooming salon, including: - Main slogan (10-15 words) - Sub-slogan (15-20 words) - Service tagline (8-12 words) - Brand promise (10-15 words) The slogans should reflect professional pet grooming and care services while emphasizing thoughtful care for pet owners. The slogans should be catchy, memorable, with warm and friendly wording, and a light, cheerful tone. Response **Main Slogan:** Pamper Your Pet, Delight Your Heart. *Sub-Slogan:** Expert grooming and loving care, making your furry friend feel fabulous, inside and out. *Service Tagline:** Gentle paws, happy pets, stress-free smiles. *Brand Promise:** Quality grooming, compassionate care, every time.
Table 16: Bad case of length requirement. Query from WritingBench with index 539, response generated by Gemini-1.5-Pro.
Figure 6: Scores of different models across various input lengths on the WritingBench.
Figure 7: Scores of different models across various output lengths on the WritingBench.